OpenStudy (anonymous):

integrate \[{2 \over 3(sec x-1)}\]

OpenStudy (anonymous):

If you could write out the steps, that'd be nice. I just keep getting the wrong answer.

OpenStudy (chaise):

Bring the 2/3 out the front of the integration and you are left with 1/sec(x-1) in the denominator with 1 in the numerator. 1/sec(x-1) = cos(x-1) integrate like normal. (make sure to bring the 2/3 out the front of the equation again) you get: -2/3sin(1-x)+c I do hope this is right. I've not yet mastered integration, but maybe this helped.

OpenStudy (anonymous):

yeah, I see what you're getting at. Thanks! let's master calc together ^^

OpenStudy (anonymous):

is this goes like this put secx = 1/cosx and then q= integrate(2/3 (cosx/(1-cosx)) =2/3((1/(1-cosx) -1 ) = 2/3 (1/2 cosec^x/2 -1) on integration 2/3 (-cotx/2 -x) + c

OpenStudy (chaise):

or maybe I am wrong..................................

OpenStudy (anonymous):

thanks all. I know what I have to do now.

OpenStudy (chaise):

Okay :) Goodluck in future

OpenStudy (anonymous):

You too :)